Hayhayz (hayhayz):

what is the relationship between exponential and logarithmic functions? what is an example of an exponential function and its inverse?

Hayhayz (hayhayz):

@Awolflover1 can you help?

OpenStudy (mayankdevnani):

they are inverse of each other

Hayhayz (hayhayz):

And an example?

OpenStudy (mayankdevnani):

\[\large \bf y=e^x\] Take log both sides, \[\large \bf \ln y=x\]

Hayhayz (hayhayz):

ok would this example be right too? b^x=a <--> logb a=x

OpenStudy (mayankdevnani):

yeah ! correct